OpenStudy (supernova_sonntag):

Question! ~I already have the answer-I am just not able to get to it.

OpenStudy (supernova_sonntag):

Find tan theta if \[\sec \theta=\frac{ \sqrt{37} }{ 6 }\] and sin theta is less than zero. I know the answer is -1/6

OpenStudy (mathmale):

You don't have to do this, but one approach would be to rewrite\[\sec \theta=\frac{ \sqrt{37} }{ 6 }\]

OpenStudy (mathmale):

as\[\cos \theta = \frac{ 6 }{\sqrt{37} }\]

OpenStudy (mathmale):

Since the cosine is definied as (adjacent side) / (hypotenuse), you now know the lengths of each. If you want to find the tangent of theta, you need to know both the opp. side and the adj. side. How would you find the opp side?

OpenStudy (supernova_sonntag):

oh I see! I've got it now, thank you. I didn't think to rewrite it like that, but after I see how the triangle works.